I have four questions
1) quadritic equation variable x having the given numbers as solutions. Type in standard equation form.
ax^2+bx+c=0 only solution is 7
2) JIm mows the lawn in 7 hours Sue mows the lawn in 4 hours. How much time would it take them to mow the lawn together?
3) Variation constant and an equation of variation where y vaires directly as x and y=6 when x=3? Variation constant k Equation variation is y
4) (20b^3+17b^2+18b+42) divided (4b+5)
Thanks in advance you guys rock!!

Answer by user_dude2008(1862)

1)

only solution: 7 ---> x=7 occurs twice


x=7 or x=7

x-7=0 or x-7=0


(x-7)(x-7)=0

x^2-14x+49=0

Ans: equation is y = x^2-14x+49


2)

Jim's rate: 1/7 lawn/hour
Sue's rate: 1/4 lawn/hour


Add rates and multiply by "t". Set equal to 1:


t(1/7+1/4)=1


t(11/28)=1


t=28/11

t=2.5455

Ans: takes about 2.5455 hours working together


3)

y=kx

6=k(3)

6/3=k

k=2

Equation: y=2x


4)


Somebody Already answered ---> http://www.algebra.com/algebra/homework/Graphs/Graphs.faq.question.195950.html